# Sequential Attention-based Sampling for Histopathological Analysis

### v2

@@ 4.6 Effect of observation budget: Calibration
We have shown that SASHA achieves near-sota performance at a fraction of the inference time, even with sampling 10% or 20% of the patches at high resolution. Is there any advantage to sampling more patches, other than achieving a marginal improvement accuracy, especially when the inference time increases almost linearly with the proportion of high-resolution patches sampled. Here, we show empirically that sampling a higher fraction of patches yields better calibrated predictions. With well calibrated models, the model’s confidence in its prediction (e.g., predictive uncertainty) is a useful indicator for when referral to a specialist is warranted [7]. We quantified the expected calibration error (ECE) [10] as a measure of the degree of calibration of the models. The ECE decreased systematically at the proportion sampled patches increased from 10% (SASHA-0.1) to 20% (SASHA-0.2) to 100% (HAFED). This pattern occurred possibly because providing the model more timesteps to sample the WSI in each RL episode enabled the model to recover from sub-optimal sampling actions in the early timesteps, and to sample the relevant portions of the WSI in the later timesteps. Surprisingly, ECE with SASHA-0.2 was lower than that of even sota attention-based models like ACMIL and DTFD (Fig. 3h). In other words, while a shorter observation budget encouraged efficient learning, a longer budget produced better calibrated predictions.
@@ 5 Conclusions and Limitations
We show that an attention-augmented deep RL agent outperforms or performs on par with sota models on WSI classification, albeit at a fraction of their memory and computational costs. Key limitations of the model include a significant impact of the observation budget on model calibration, and the need to train the model with all high resolution patches, which significantly increases training time; other limitations are discussed in Appendix A.6. Nevertheless, we demonstrate explainable, reliable, and calibrated predictions, which could enhance clinicians’ trust in model predictions.
@@ A.1 Dataset details
The CAMELYON16 dataset comprises 399 whole-slide images (WSIs) of lymph node sections. After pre-processing and segmentation, we generated non-overlapping patches of size $256\times 256$ (×$\times 3$ RGB channels). On average, this resulted in, on average, $\sim$800 patches per WSI at a scanning resolution of 5×5$\times$5 ×, and $\sim$12,500 patches at a high resolution of 20×20$\times$20 ×. We used the official test set for evaluation [3]. The training set was further divided to create a validation set (10%) while maintaining a similar proportion of normal and tumor slides in each subset.
The TCGA-NSCLC dataset comprises 1,054 whole-slide images (WSIs) obtained from the TCGA-NSCLC repository under the TCGA-LUSC and TCGA-LUAD projects. Ten low-quality slides without magnification information were discarded, and three slides with errors were ignored. We followed the same test split as in DSMIL [17]. The test set comprised a total of 213 slides: 104 from TCGA-LUAD and 109 from TCGA-LUSC. Following preprocessing, we obtained, on average, $\sim$212 patches per WSI at 5×5$\times$5 × magnification and $\sim$3,393 patches at 20×20$\times$20 × magnification. The training set was further divided to create a validation (20%) set while maintaining a similar proportion of labels in each subset.
@@ A.2.1 Patch feature distillation with HAFED
The first step of SASHA involves feature extraction and aggregation of high-resolution patch features, which specifies the RL agent’s state $S(t)$. We use a Vision Transformer (ViT) pretrained on histopathological images to extract features from each zoomed in, high-resolution patch [16]. This ViT was used as is without any additional fine tuning, and its output – a 384 dimensional embedding per patch – was provided as input to a hierarchical attention-based feature aggregator and distiller (HAFED).
@@
Table 4: Architecture of the HAFED model. Layers 1-3 corresponds to feature aggregator (FA) and Layers 4-7 corresponds to classifier. $N$ ($\sim$800 for camelyon-16) is the number of patches in low resolution WSI, $k$ (16) is number of sub-patches in high-resolution per one patch in low-resolution, $d$ (384) feature embedding dimension of one patch. $U_{i}\in\mathbb{R}^{16\times 384}$, $A_{i}\in\mathbb{R}^{16\times 128}$, $B_{i}\in\mathbb{R}^{16\times 128}$, $C\in\mathbb{R}^{N\times 128}$, $D\in\mathbb{R}^{N\times 128}$, $h\in\mathbb{R}^{384}$.
@@
As described in Section 3.2, the first stage of HAFED acts as a feature aggregator (FA); the architecture of this network is shown in Table 4. We employed a single attention branch, which takes high resolution features $U\in\mathbb{R}^{N\times k\times d}$ and assigns attention scores for each of the $k$ sub patches. The inner product between the attention scores and $U$ produces $V\in\mathbb{R}^{N\times d}$, which is in same dimensionality as the low-resolution features ($Z\in\mathbb{R}^{N\times d}$). The second stage of HAFED acts as a classifier, which is trained end-to-end with the feature aggregator. Here, we use five attention branches for the CAMELYON16 dataset and one for the TCGA-NSCLC dataset; the number of branches was determined as a hyperparameter based on the respective validation set (Table 6). This stage takes $V$ as input, assigns attention scores across the N patches, and computes the inner product between the scores and $V$ to produce a 384-dimensional embedding representing the entire WSI. The architecture of this network is shown in Fig. 2a. We applied masking to the top four and top eight patches with a probability of 0.4 in the first and second stages, respectively (Table 6).
@@
HAFED is trained with an annealed learning rate schedule starting at 4 ×$\times 10^{-4}$, with the AdamW optimizer and a weight decay of 1 ×$\times 10^{-4}$. The HAFED module is trained end-to-end using the following loss components: (1) a cross-entropy loss between the final predicted label and the ground truth label, (2) a cross-entropy loss applied independently to the predictions from each attention branch in second stage and ground truth label, and (3) a dissimilarity loss that encourages diversity among the attention scores of the different branches in the second stage. This dissimilarity loss encourages each branch to focus on distinct tumor features, thereby improving the model’s ability to capture complementary diagnostic features[35].
@@
During inference, at each time step, the first stage of HAFED (FA), receives as input $k$ zoomed-in sub-patch images corresponding to the selected action. It assigns attention scores to estimate the relative importance of each sub-patch and aggregates their features by computing a weighted sum based on these scores; this is then used to update the state (see next subsection, Section A.2.2). In the final time step, the second stage of HAFED (classifier), takes the state representation $S_{T}$ and predicts the label $\hat{y}$. In addition to the final classification, the second stage is also used to assign rewards during SASHA’ RL agent training (see subsection Section A.2.3).
A.2.2 Updating WSI state with TSU
The second step of SASHA involves updating the RL agent’s state with the high-resolution patch representation extracted with HAFED. During training, the states of similar patches are concurrently updated with a targeted state update (TSU) rule. Specifically, this included all patches whose low-resolution representation’s similarity to the currently selected patch, exceeded a threshold value ($\tau$), based on cosine similarity; $\tau$ was tuned as a hyperparameter (Table 6)
The architecture of this TSU network is shown in Table 5. For each patch $a_{\tau}$ that crosses the similarity threshold, the TSU model takes as input a concatenated vector comprising (i) its low-resolution features $Z_{a_{\tau}}\in\mathbb{R}^{d}$, (ii) the low-resolution feature representation of the selected patch $a_{t}$, $Z_{a_{t}}\in\mathbb{R}^{d}$ and (iii) the high-resolution feature representation of the selected patch $V_{a_{t}}\in\mathbb{R}^{d}$. Note that while (i-ii) are obtained directly from pretrained ViT, (iii) is obtained from the FA output of the HAFED model. Concatenating these 3 features results in a $3d$-dimensional input (3×384=1152)3$\times$ 384=1152)3 × 384 = 1152 ), based on which the model predicts a $d$-dimensional high-resolution feature of the patch $a_{\tau}$, $\hat{V}_{a_{\tau}}$. The state of the selected patch alone is updated by directly copying its high-resolution HAFED (FA) output $V_{a_{t}}$, into its (corresponding) entry in the WSI state $S(t)$. In this manner, the TSU integrates information from the low-resolution representation of patches similar to the currently visited patch, with the high-resolution representation of visited patches to update the WSI state.
During training the TSU network is trained by sampling patches randomly, and updating their states based on the predicted high-resolution feature of each patch ($\hat{V}$) and its actual high-resolution feature ($V$) obtained from HAFED’s feature aggregator. Once a patch state is updated – either directly, by visiting it, or indirectly, by the similarity-based rule – its index is masked out, and it is no longer visited during training. The TSU model is trained for 500 epochs with an Adam optimizer and a learning rate of 1 ×$\times 10^{-4}$. The best model was chosen based on the minimum of the validation loss.
Layer Input Dim Output Dim # Parameters
FC-1 1152 768 885,504
FC-2 768 384 295,296
Layer Norm - 384 384
Table 5: Architecture of the TSU model. The input is a $11523d=$ 1152 dimensional vector ($d=384$; see text for details). The final output is $d$-dimensional. FC: fully connected layer with bias.
A.2.3 Actor and Critic
The final component of SASHA is a reinforcement learning (RL) agent responsible for sampling, sequentially, the most relevant patches for high-resolution analysis. This selection is formulated as a sequential decision-making problem, where the agent learns to identify diagnostically informative regions within each slide. We adopt the Actor-Critic framework and train the agent using the Proximal Policy Optimization (PPO) algorithm, a widely used on-policy method known for its stability and effectiveness in complex RL tasks [28].
The purpose of the actor network is to learn a policy – a conditional probability distribution over actions given the current state $S_{t}$. In our setup, the environment is described as, the state is initialized with low-resolution features ($Z\in\mathbb{R}^{N\times d}$), the action is selected from a discrete action space $A\in\{1,2,\ldots,N\}$ (indices of low-resolution patches), and the reward function is specified as the negative of binary cross-entropy (BCE) loss between the predicted WSI label, at time step $t,$ and the ground truth. Given the definitions of state, action and reward, the actor and critic networks are designed as follows.
The Actor network follows the same architecture as Layers 1–3 in Table 4. A softmax is applied to the output attention scores to obtain a categorical distribution (i.e., the policy) over the $N$ patches, from which an action $a_{t}$ (i.e., a patch index) is sampled during training. To reduce redundancy, once a patch is selected, its index is masked in all subsequent time steps – by setting its corresponding logit to negative infinity – effectively removing it from consideration for future visits. the rationale behind this policy is that zooming into the same patch multiple times will not yield additional information, as the patch content remains unchanged. After selecting $a_{t}$, the Target State Updater (TSU) module is invoked to update the current state $S_{t}$ to the next state $S_{t+1}$, thereby incorporating the high-resolution information from the selected patch into the state representation.
The Critic network also mirrors the backbone of actor network i.e architecture of Layers 1–3 in Table 4, with an additional final layer that takes the aggregated state representation ($\alpha\cdot S_{t}$), and predicts the reward-to-go or expected return given the current state $S_{t}$. The reward-to-go is used in calculating the generalized advantage estimate (GAE, [27]). During training the actual reward-to-go from the given state is known, and the critic network is trained by minimizing MSE loss between the predicted and actual value.
[…]
During inference, given the current state $S_{t}$, the Actor network outputs a policy in the form of a categorical probability distribution over the $N$ patches. The action $a_{t}$ is then selected deterministically as the patch index with the highest probability. In a later section (Section A.3.2) we explore the effect of alternative, stochastic selection policies during inference.

### v3

@@ 4.6 Effect of observation budget: Calibration
We have shown that SASHA achieves near-sota performance at a fraction of the inference time, even with sampling 10% or 20% of the patches at high resolution. Is there any advantage to sampling more patches, other than achieving a marginal improvement accuracy, especially when the inference time increases almost linearly with the proportion of high-resolution patches sampled. Here, we show empirically that sampling a higher fraction of patches yields better calibrated predictions. With well calibrated models, the model’s confidence in its prediction (e.g., predictive uncertainty) is a useful indicator for when referral to a specialist is warranted [Dj_and_jim]. We quantified the expected calibration error (ECE) [guo2017calibration] as a measure of the degree of calibration of the models. The ECE decreased systematically at the proportion sampled patches increased from 10% (SASHA-0.1) to 20% (SASHA-0.2) to 100% (HAFED). This pattern occurred possibly because providing the model more timesteps to sample the WSI in each RL episode enabled the model to recover from sub-optimal sampling actions in the early timesteps, and to sample the relevant portions of the WSI in the later timesteps. Surprisingly, ECE with SASHA-0.2 was lower than that of even sota attention-based models like ACMIL and DTFD (Fig.˜3h). In other words, while a shorter observation budget encouraged efficient learning, a longer budget produced better calibrated predictions.
@@ 5 Conclusions and Limitations
We show that an attention-augmented deep RL agent outperforms or performs on par with sota models on WSI classification, albeit at a fraction of their memory and computational costs. Key limitations of the model include a significant impact of the observation budget on model calibration, and the need to train the model with all high resolution patches, which significantly increases training time; other limitations are discussed in Appendix A.6. Nonetheless, we demonstrate explainable, reliable, and calibrated predictions, which could enhance clinicians’ trust in model predictions.
Acknowledgements
This research was supported by a Department of Science and Technology SwarnaJayanti Fellowship, a Google Research Grant, and a Pratiksha Trust Intramural Grant (all to DS), Ministry of Education Scholarships (to TG and NM), an Institute of Eminence (IoE) Grant (to IISc), a Walmart Centre for Tech Excellence at IISc, the Indo–French Centre for the Promotion of Advanced Research Grant (CEFIPRA; Project 7102–1), Kotak IISc FinTech Grant, a DST–SERB Core Research Grant (CRG/2021/008330), and a Pratiksha Trust Young Investigator Award (all to GT). We are very grateful to Karm Patel, Mainak Biswas, Aditi Saxena, and Tej Pratap Yadav for help with model development and simulations, and Raj Jain for help with the original draft of the manuscript. The authors declare no competing interests.
@@ A.1 Dataset details
The CAMELYON16 dataset comprises 399 whole-slide images (WSIs) of lymph node sections. After pre-processing and segmentation, we generated non-overlapping patches of size $256\times 256$ ($\times 3$ RGB channels). On average, this resulted in, on average, $\sim$800 patches per WSI at a scanning resolution of $5\times$, and $\sim$12,500 patches at a high resolution of $20\times$. We used the official test set for evaluation [bejnordi2017diagnostic]. The training set was further divided to create a validation set (10%) while maintaining a similar proportion of normal and tumor slides in each subset.
The TCGA-NSCLC dataset comprises 1,054 whole-slide images (WSIs) obtained from the TCGA-NSCLC repository under the TCGA-LUSC and TCGA-LUAD projects. Ten low-quality slides without magnification information were discarded, and three slides with errors were ignored. We followed the same test split as in DSMIL [dsmil2021]. The test set comprised a total of 213 slides: 104 from TCGA-LUAD and 109 from TCGA-LUSC. Following preprocessing, we obtained, on average, $\sim$212 patches per WSI at $5\times$ magnification and $\sim$3,393 patches at $20\times$ magnification. The training set was further divided to create a validation (20%) set while maintaining a similar proportion of labels in each subset.
@@
The CamelyonPlus dataset[ling2025camelyonplus], comprises 1,347 whole‐slide images (WSIs) derived from the curated CAMELYON16 and CAMELYON17 collections, after removal of low-quality slides and re-labeling. It involves a four-class classification. We created training, validation, and test splits to ensure class balance across all four labels. A summary of the splits and labels can be found in Table˜4.
Dataset Train Val Test
Neg. Mi-m. Ma-m. ITC Neg. Mi-m. Ma-m. ITC Neg. Mi-m. Ma-m. ITC
Camelyon+ 608 122 174 36 130 26 38 9 131 26 38 9
Table 4: Summary of Camelyon+ dataset. The WSIs have the following 4-class slide-level labels: Negative (Neg.), Micro-metastasis (Mi-m.), Macro-metastasis (Ma-m.), and Isolated Tumor Cells (ITC). The table shows the number of WSIs for each class in the training, validation, and test sets.
The BRACS (BReAst Carcinoma Subtyping) dataset comprises 547 labeled WSIs obtained from the public repository [brancati2022bracs]. This dataset involves a 7-class classification task, characterizing each tumor as one of three benign (BT), two atypical (AT), or two malignant (MT) sub-types. We ignored the “normal” subtype in the benign (BT) class, and each set of subtypes within BT, AT, and MT was grouped into its respective class. Additionally, a small number of slides (7/503) that lacked all relevant levels of magnification were excluded. After preprocessing, we use the official dataset split across train, validation, and test split with details outlined in Table˜5.
Benign (BT) Atypical (AT) Malignant (MT)
Train 171 52 135
Val 16 14 21
Test 25 23 16
Table 5: BRACS dataset distribution for 3-class classification (Benign, Atypical, Malignant) across training, validation, and test splits.
@@ A.2.1 Patch feature distillation with HAFED
The first step of SASHA involves feature extraction and aggregation of high-resolution patch features, which specifies the RL agent’s state $S(t)$. We use a Vision Transformer (ViT) pretrained on histopathological images to extract features from each zoomed in, high-resolution patch [kang2023ssl]. This ViT was used as is without any additional fine tuning, and its output – a 384 dimensional embedding per patch – was provided as input to a hierarchical attention-based feature aggregator and distiller (HAFED).
@@
Table 6: Architecture of the HAFED model. Layers 1-3 correspond to the feature aggregator (FA) and layers 4-7 correspond to the classifier. NN ($\sim$800 for camelyon-16) is the number of patches in low-resolution WSI, $k$ (16) is the number of sub-patches in high-resolution per one patch in low-resolution, $d$ (384) feature embedding dimension of one patch. $U_{i}\in\mathbb{R}^{16\times 384}$, $A_{i}\in\mathbb{R}^{16\times 128}$, $B_{i}\in\mathbb{R}^{16\times 128}$, $C\in\mathbb{R}^{N\times 128}$, $D\in\mathbb{R}^{N\times 128}$, $h\in\mathbb{R}^{384}$.
@@
As described in Section˜3.2, the first stage of HAFED acts as a feature aggregator (FA); the architecture of this network is shown in Table˜6. We employed a single attention branch, which takes high resolution features $U\in\mathbb{R}^{N\times k\times d}$ and assigns attention scores for each of the $k$ sub patches. The inner product between the attention scores and $U$ produces $V\in\mathbb{R}^{N\times d}$, which is in same dimensionality as the low-resolution features ($Z\in\mathbb{R}^{N\times d}$). The second stage of HAFED acts as a classifier, which is trained end-to-end with the feature aggregator. Here, we use five attention branches for the CAMELYON16 dataset and one for the TCGA-NSCLC dataset; the number of branches was determined as a hyperparameter based on the respective validation set (Table˜7). This stage takes $V$ as input, assigns attention scores across the N patches, and computes the inner product between the scores and $V$ to produce a 384-dimensional embedding representing the entire WSI. The architecture of this network is shown in Fig.˜2a. We applied masking to the top four and top eight patches with a probability of 0.4 in the first and second stages, respectively (Table˜7).
@@
HAFED is trained with an annealed learning rate schedule starting at 4 $\times 10^{-4}$, with the AdamW optimizer and a weight decay of 1 $\times 10^{-4}$. The HAFED module is trained end-to-end using the following loss components: (1) a cross-entropy loss between the final predicted label and the ground truth label, (2) a cross-entropy loss applied independently to the predictions from each attention branch in second stage and ground truth label, and (3) a dissimilarity loss that encourages diversity among the attention scores of the different branches in the second stage. This dissimilarity loss encourages each branch to focus on distinct tumor features, thereby improving the model’s ability to capture complementary diagnostic features[zhang2023acmil].
@@
During inference, at each time step, the first stage of HAFED (FA), receives as input $k$ zoomed-in sub-patch images corresponding to the selected action. It assigns attention scores to estimate the relative importance of each sub-patch and aggregates their features by computing a weighted sum based on these scores; this is then used to update the state (see next subsection, Section˜A.2.2). In the final time step, the second stage of HAFED (classifier), takes the state representation $S_{T}$ and predicts the label $\hat{y}$. In addition to the final classification, the second stage is also used to assign rewards during SASHA’ RL agent training (see subsection Section˜A.2.3).
@@
A.2.2 Updating WSI state with TSU
The second step of SASHA involves updating the RL agent’s state with the high-resolution patch representation extracted with HAFED. During training, the states of similar patches are concurrently updated with a targeted state update (TSU) rule. Specifically, this included all patches whose low-resolution representation’s similarity to the currently selected patch, exceeded a threshold value ($\tau$), based on cosine similarity; $\tau$ was tuned as a hyperparameter (Table˜7)
The architecture of this TSU network is shown in Table˜8. For each patch $a_{\tau}$ that crosses the similarity threshold, the TSU model takes as input a concatenated vector comprising (i) its low-resolution features $Z_{a_{\tau}}\in\mathbb{R}^{d}$, (ii) the low-resolution feature representation of the selected patch $a_{t}$, $Z_{a_{t}}\in\mathbb{R}^{d}$ and (iii) the high-resolution feature representation of the selected patch $V_{a_{t}}\in\mathbb{R}^{d}$. Note that while (i-ii) are obtained directly from pretrained ViT, (iii) is obtained from the FA output of the HAFED model. Concatenating these 3 features results in a $3d$-dimensional input ($3\times 384=1152)$, based on which the model predicts a $d$-dimensional high-resolution feature of the patch $a_{\tau}$, $\hat{V}_{a_{\tau}}$. The state of the selected patch alone is updated by directly copying its high-resolution HAFED (FA) output $V_{a_{t}}$, into its (corresponding) entry in the WSI state $S(t)$. In this manner, the TSU integrates information from the low-resolution representation of patches similar to the currently visited patch, with the high-resolution representation of visited patches to update the WSI state.
During training the TSU network is trained by sampling patches randomly, and updating their states based on the predicted high-resolution feature of each patch ($\hat{V}$) and its actual high-resolution feature ($V$) obtained from HAFED’s feature aggregator. Once a patch state is updated – either directly, by visiting it, or indirectly, by the similarity-based rule – its index is masked out, and it is no longer visited during training. The TSU model is trained for 500 epochs with an Adam optimizer and a learning rate of 1 $\times 10^{-4}$. The best model was chosen based on the minimum of the validation loss.
Layer Input Dim Output Dim # Parameters
FC-1 1152 768 885,504
FC-2 768 384 295,296
Layer Norm - 384 384
Table 8: Architecture of the TSU model. The input is a $3d=1152$ dimensional vector ($d=384$; see text for details). The final output is $d$-dimensional. FC: fully connected layer with bias.
A.2.3 Actor and Critic
The final component of SASHA is a reinforcement learning (RL) agent responsible for sampling, sequentially, the most relevant patches for high-resolution analysis. This selection is formulated as a sequential decision-making problem, where the agent learns to identify diagnostically informative regions within each slide. We adopt the Actor-Critic framework and train the agent using the Proximal Policy Optimization (PPO) algorithm, a widely used on-policy method known for its stability and effectiveness in complex RL tasks [schulman2017ppo].
The purpose of the actor network is to learn a policy – a conditional probability distribution over actions given the current state $S_{t}$. In our setup, the environment is described as, the state is initialized with low-resolution features ($Z\in\mathbb{R}^{N\times d}$), the action is selected from a discrete action space $A\in\{1,2,\ldots,N\}$ (indices of low-resolution patches), and the reward function is specified as the negative of cross-entropy (CE) loss between the predicted WSI label, at time step $t,$ and the ground truth. Given the definitions of state, action and reward, the actor and critic networks are designed as follows.
The Actor network follows the same architecture as Layers 1–3 in Table˜6. A softmax is applied to the output attention scores to obtain a categorical distribution (i.e., the policy) over the $N$ patches, from which an action $a_{t}$ (i.e., a patch index) is sampled during training. To reduce redundancy, once a patch is selected, its index is masked in all subsequent time steps – by setting its corresponding logit to negative infinity – effectively removing it from consideration for future visits. the rationale behind this policy is that zooming into the same patch multiple times will not yield additional information, as the patch content remains unchanged. After selecting $a_{t}$, the Target State Updater (TSU) module is invoked to update the current state $S_{t}$ to the next state $S_{t+1}$, thereby incorporating the high-resolution information from the selected patch into the state representation.
The Critic network also mirrors the backbone of actor network i.e architecture of Layers 1–3 in Table˜6, with an additional final layer that takes the aggregated state representation ($\alpha\cdot S_{t}$), and predicts the reward-to-go or expected return given the current state $S_{t}$. The reward-to-go is used in calculating the generalized advantage estimate (GAE, [gae]). During training the actual reward-to-go from the given state is known, and the critic network is trained by minimizing MSE loss between the predicted and actual value.
Both networks are trained for 15 epochs using the AdamW optimizer with an annealing learning rate starting from $1\times 10^{-5}$ and a weight decay of $1\times 10^{-3}$. Network hyperparameters are reported in Table˜7. During training of the Actor and Critic networks, all other components of SASHA are frozen (i.e., HAFED, TSU), and only the actor and critic networks are updated. We select the best policy across epochs as the one that minimizes the validation loss while simultaneously maximizing the sum of the validation AUROC and F1 scores ($\mathrm{min}_{\pi}[\mathrm{Loss+2-(AUROC+F1)}]$). This formulation ensures that the selected policy balances both accuracy (AUROC and F1) and calibration (validation loss).
During inference, given the current state $S_{t}$, the Actor network outputs a policy in the form of a categorical probability distribution over the $N$ patches. The action $a_{t}$ is then selected deterministically as the patch index with the highest probability. In a later section (Section˜A.3.3) we explore the effect of alternative, stochastic selection policies during inference.